Given a situation where the corporate tax rate is 34%, and the personal

tax rate on dividends is 28%, what must the personal tax rate on interest

be to achieve the Miller Model equilibrium (i.e. where V = Vu)?
Answer #1

Asnwer:

GIven

Corporate tax Tc=34%

Dividend tax Tp =28%

Let Td be the tax on interest

(1-Td/(1-Tc)=(1-Tp)

1-Td=(1-28%)*(1-24%)

Td=52.48%